OpenStudy (anonymous):

please answer quickly : Solve 5 (^x + 5) = 9x. *logs*

OpenStudy (anonymous):

See the other problem's solution and do it same way.

OpenStudy (anonymous):

x ≈ –3.63 x ≈ 13.69 x ≈ 0.07 x ≈ –0.28 i think its the first one!! is that right?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Are you sure the other side is 9x and not 9^x?

OpenStudy (anonymous):

your right! 9^x

OpenStudy (anonymous):

is it the first option?..... or am i off

OpenStudy (anonymous):

(x+5) * ln(5) = x * ln(9) x (ln(5) - ln(9)) = -5*ln(5) x = -5(ln(5))/(ln(5)-ln(9))

OpenStudy (anonymous):

13.69

OpenStudy (anonymous):

thanks:) i posted one last question, pleas help for a medal

OpenStudy (anonymous):

Sounds good. I don't have a scientific calculator to truly calculate and confirm.

OpenStudy (paxpolaris):

yes, it is. you can just google -5(ln(5))/(ln(5)-ln(9))